# Distillation of multipartite entangled states for arbitrary subsets of parties in noisy quantum networks of increasing size

### v1

@@ Abstract
Quantum network states are multipartite states built from distributing pairwise entanglement among parties and underpin the paradigm of quantum networks for quantum information processing. In this work we introduce the problem of partial distillability in noisy quantum networks. This corresponds to the possibility of creating locally starting from a single copy of a quantum network state with mixed entangled links supporting a constant amount of noise an arbitrary pure state for an arbitrary subset of parties with fidelity as close to 1 as desired as the size of the network increases. While we prove an obstruction to multipartite distillation protocols after teleportation with channels with constant noise, we show that partial distillability is indeed possible if certain well-established graph-theoretic parameters that measure the connectivity in the network grow fast enough with its size. We give necessary as well as sufficient conditions for partial distillability in terms of these parameters and we moreover provide explicit constructions of networks with partial distillability and a relatively slow connectivity growth.
@@ II.1 Graph theory
All graphs in this article are finite and simple. Thus, a graph is a pair $G=(V,E)$, where $V$ is a finite set of vertices and $E$ a set of edges, consisting of unordered pairs $(u,v)$ of vertices $u,v\in V$ ($u\neq v$). The order of $G$, denoted $|G|$, is the cardinality of the set of vertices $V$ and the size of $G$ is the cardinality of the set of edges $E$. Two vertices $u,v\in V$ are said to be adjacent if $(u,v)\in E$. We denote the set of edges adjacent to a vertex $v\in V$ by $E(v)$. The degree of a vertex $v\in V$, $\delta(v)=|E(v)|$, is number of adjacent vertices to it, and we denote by $\delta_{\mathrm{min}}(G)=\min_{v\in V}\delta(v)$ ($\delta_{\mathrm{max}}(G)=\max_{v\in V}\delta(v)$) the minimal (maximal) degree of a graph $G$. A path is a sequence of edges joining subsequent vertices, $(v_{0},v_{1}),(v_{1},v_{2}),\dots,(v_{\ell-1},v_{\ell})$, and when $v_{j}\neq v_{i}$ for $i\neq j$ the path is said to be simple. A cycle is a path starting and ending on the same vertex. The length of a path is the number of edges in the path and the distance between two vertices $u,v\in V$, $d_{G}(u,v)$, is the length of the shortest path joining $u$ and $v$.
@@ II.2 Quantum (network) states and distillability
In this article we will always consider finite-dimensional Hilbert spaces. A quantum state is given by any positive semidefinite unit-trace operator $\rho$, referred to as density operator, acting on a given Hilbert space $\mathcal{H}$. The set of all density operators on $\mathcal{H}$ is denoted by $D(\mathcal{H})$. When a density operator is a rank-1 projector we say that the corresponding state is pure. When a quantum system is composed by $N$ subsystems, labelled by the elements of $[N]:=\{1,2,\ldots,N\}$, the corresponding Hilbert space is given by $\mathcal{H}=\bigotimes_{k=1}^{N}\mathcal{H}_{k}$, where $\mathcal{H}_{k}$ is the Hilbert space corresponding to party $k$. In the bipartite case, $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$, we denote by
@@
$\rho(p)=p\phi^{+}+\frac{(1-p)}{d^{2}}\mbox{$1{\bf l}$},$ (2)
where $p$ is referred to as the visibility and $1{\bf l}$ stands for the identity operator on $\mathcal{H}$. It is known that the isotropic state is entangled if and only if (iff) $p>1/(d+1)$ [24] and it is a common noise model in which one produces a maximally entangled state with probability $p$ and white noise with probability $1-p$.
@@
An $N$-partite quantum network state is described by a graph $G=(V,E)$ such that $|G|=N$, where the elements of $V$ represent the parties and the elements of $E$ that the parties corresponding to the vertices in a given edge share some bipartite state to be specified. Thus, we denote the total Hilbert space by $\mathcal{H}_{G}$ and $\mathcal{H}_{G}=\bigotimes_{v\in V}\mathcal{H}_{v}$, where $\mathcal{H}_{v}$ represents the Hilbert space corresponding to the party represented by the vertex $v\in V$. Notice that there is a bijective identification between the elements of $V$ and those of $[N]$ and, thus, in a slight abuse of notation, we will refer to both vertices and parties as elements of these two sets equally. If vertices $u$ and $v$ share the state $\rho$, then we write $\rho_{e}$, where $e=(u,v)\in E$, and $\rho_{e}\in D(\mathcal{H}_{e})$, where $\mathcal{H}_{e}=\mathcal{H}_{ue}\otimes\mathcal{H}_{ve}$. Thus, $\mathcal{H}_{v}=\bigotimes_{e\in E(v)}\mathcal{H}_{ve}$. Furthermore, it also holds that $\mathcal{H}_{G}=\bigotimes_{e\in E}\mathcal{H}_{e}$, even though $\mathcal{H}_{e}$ does not correspond to the local Hilbert space of any party.
@@ II.3 Noisy teleportation
The main ingredient in the LOCC protocols that we are going to consider is teleportation [28, 29]. We denote by $\mathcal{T}:D(\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3})\to D(\mathcal{H}_{3})$ the channel implementing the standard teleportation protocol for which
$\mathcal{T}(\rho\otimes\phi^{+})=\rho$ (7)
for all $\rho\in D(\mathcal{H}_{1})$. We will be often using the entangled links in an isotropic network state to teleport quantum states through the network, which will be noisy when the visibility satisfies $p<1$. Applying the same protocol by using an isotropic state instead of a maximally entangled state yields
$\mathcal{T}(\rho\otimes\rho(p))=p\rho+(1-p)\frac{\mbox{$1{\bf l}$}}{d}.$ (8)
We define the noisy teleportation channel $\mathcal{T}_{p}:D(\mathcal{H}_{1})\to D(\mathcal{H}_{3})$ as
$\mathcal{T}_{p}(\rho):=\mathcal{T}(\rho\otimes\rho(p)),$ (9)
@@
$\mathcal{T}_{p}(|i\rangle\langle j|)=p|i\rangle\langle j|+(1-p)\delta_{ij}\frac{\mbox{$1{\bf l}$}}{d}.$ (10)
@@
Another option is to directly distribute a multipartite quantum state to some set of parties $V$ by means of a star graph $S=(V^{\ast},E)$, where $V^{*}=V\cup\{v_{0}\}$ and $v_{0}$ is its center. This additional party can locally prepare any state $\rho\in D(\mathcal{H}^{\prime}_{V})$ and teleport each share of it to each of the parties in $V$ by means of each edge in the star graph using the corresponding isotropic state as a resource for noisy teleportation. Thus, this corresponds to implementing the LOCC map from $D(\mathcal{H}_{S})$ to $D(\mathcal{H}^{\prime}_{V})$ given by $(\mathcal{T}_{p})^{\otimes|V|}(\rho)$ for any chosen $\rho\in D(\mathcal{H}^{\prime}_{V})$. If the parties in $V$ are linked to many additional parties giving rise to star graphs with different centers, they can iterate this process in order to end up with many noisy copies of the target state, to which they can apply multipartite distillation protocols as a final stage. We will in particular later on consider the paradigmatic GHZ $n$-qubit state
@@
$\sigma_{k}^{C}=\frac{1}{2}\left(\bigotimes_{\ell\notin C}|0\rangle_{\ell}\langle 0|+\bigotimes_{\ell\notin C}|1\rangle_{\ell}\langle 1|\right)\otimes\frac{\mbox{$1{\bf l}$}_{C}}{2^{k}}.$ (15)
@@ III Limitation of multipartite distillation protocols
where #$\#(j,k)$ is the number of elements in $\{C\in\mathcal{C}(n,k)\mid\langle\Psi_{j}^{\pm}|\sigma_{k}^{C}|\Psi_{j}^{\pm}\rangle\neq 0\}$; that is, the number of non-vanishing terms in the sum $\sum_{C\in\mathcal{C}(n,k)}\langle\Psi_{j}^{\pm}|\sigma_{k}^{C}|\Psi_{j}^{\pm}\rangle$. Let $C\in\mathcal{C}(n,k)$, $j_{1}j_{2}\dots j_{n-1}$ the binary expression of $j$ and $j_{n}=0$ in order to use the same notation as above. Given $j$, Lemma 3 tells us #$\#(j,k)$ is the number of elements $C\in\mathcal{C}(n,k)$ such that for every $\ell_{1},\ell_{2}\notin C$ it holds $j_{\ell_{1}}=j_{\ell_{2}}$. For $i=0,1$, let us denote by #$\#_{i}(j,k)$ the number of elements $C\in\mathcal{C}(n,k)$ such that $j_{\ell}=i$ implies $\ell\in C$. By definition, it holds #$\#(j,k)=\#_{1}(j,k)+\#_{0}(j,k)$. It follows now that #$\#_{1}(j,k)=0$ whenever $k<w_{H}(j_{1}j_{2}\dots j_{n-1}0)=w_{H}(j)$. Otherwise, #$\#_{1}(j,k)$ must be the number of different subsets of $[n]$, with size $k$, containing the positions in which $j_{1}j_{2}\dots j_{n-1}0$ has ones. Since these $w_{H}(j)$ positions have to be assigned to $C$ always, #$\#_{1}(j,k)$ is given by the different ways in which the remaining $n-w_{H}(j)$ positions can fill the remaining $k-w_{H}(j)$ slots in $C$, i.e.
#$\#_{1}(j,k)=\binom{n-w_{H}(j)}{k-w_{H}(j)}.$ (25)
On the other hand, for #$\#_{0}(j,k)$ the $n-w_{H}(j)$ positions where $j_{1}j_{2}\dots j_{n-1}0$ has a zero must always be assigned to $C$ and, therefore, #$\#_{0}(j,k)$ is given by the different ways in which the remaining $w_{H}(j)$ positions can fill the remaining $k-(n-w_{H}(j))$ slots in $C$, i.e.
#$\#_{0}(j,k)=\binom{w_{H}(j)}{k+w_{H}(j)-n}.$ (26)
@@ IV.1 A sufficient condition for partial distillability
If $(G_{n})_{n\in\mathbb{N}}$ is graph sequence such that $\delta_{\mathrm{min}}(G_{n})$ is $\Omega(|G_{n}|)$ and $\lambda(G_{n})$ is $\omega(1)$ (i.e. $\lim_{n\to\infty}\lambda(G_{n})=\infty$), then it has partial distillability.
@@
and we assume that $M\geq 1$ since otherwise the claim is trivially true. For each $1\leq m\leq M$, we will construct the spider graph $S_{m}$ out of $|V_{0}|-1$ paths $P_{m,k}$, $11\leq k\leq|V_{0}|-$ 1 joining $v$ with $v_{k}$. To find these paths we use an inductive construction based on a family of subgraphs $G_{m,k}$, with $G_{1,1}=G$. If $G_{m,k}$ is connected, we define $P_{m,k}$ to be the shortest path inside $G_{m,k}$ joining $v$ with $v_{k}$. We then define $G_{m,k+1}$ to be the subgraph of $G_{m,k}$ obtained by removing all the edges in $P_{m,k}$. Once the last leg of the spider is obtained (i.e., the path $P_{m,|V_{0}|-1}$), we define $S_{m}$ to be the union of every path in $\{P_{m,k}\}_{k=1}^{|V_{0}|-1}$ and start building the next spider graph, starting over from $G_{m+1,1}=G_{m,|V_{0}|}$. Notice then that by construction the spider graphs are guaranteed to be edge-disjoint.
For the previous construction to work and to prove the claim, we need to ensure for all $1\leq m\leq M$ and all $11\leq k\leq|V_{0}|-$ 1 that $G_{m,k}$ is connected and that the length of the path $P_{m,k}$ is upper bounded by $5/c$. We know that $G_{1,1}=G$ is connected by hypothesis, and, by Eq. (34), that $\operatorname{diam}(G_{1,1})\leq 3/c<5/c$, so that $P_{1,1}$ fulfills the desired bound on its length. Thus, we can use induction over $j$, where we use the bijective mapping $j(k,m):=(m-1)|V_{0}|+k$, assuming that $11\leq j\leq M|V_{0}|-$ 1 (so that $1\leq m\leq M$ and $11\leq k\leq|V_{0}|-$ 1). We have just verified that the first element in this set satisfies our claim, so we now assume that $G_{m,k}$ is connected and that the length of $P_{m,k}$ is upper bounded by $5/c$ and we need to show that the same holds for the graph indexed by $j(k,m)+1$ as long as $j\leq M|V_{0}|-2$. Note that $G_{m,k}$ is built by removing a simple path from $G_{m,k-1}$ (or $G_{m-1,|V_{0}|-1}$ if $k=1$), and therefore in each step the minimal degree of $G_{m,k}$ can at most be decreased by 2 from the previous step, which implies
@@
We will explicitly build a protocol that achieves partial distillation under the given premises. Since $\delta_{\mathrm{min}}(G)\leq|G|-1$ for any graph $G$, $\delta_{\mathrm{min}}(G_{n})$ being $\Omega(|G_{n}|)$ implies that there exists $10<c<$ 1 and $n_{1}>0$ such that $\delta_{\mathrm{min}}(G_{n})\geq c|G_{n}|$ for every $n>n_{1}$. In the remainder of the proof we assume that $n$ is large enough so that the above inequality holds. Fix

### v2

@@ Abstract
Quantum network states are multipartite states built from distributing pairwise entanglement among parties and underpin the paradigm of quantum networks for quantum information processing. In this work we introduce the problem of partial distillability in noisy quantum networks. This corresponds to the possibility of distilling by local operations and classical communication an arbitrary pure state for an arbitrary subset of parties starting from a single copy of a quantum network state of an increasing number of parties that share noisy (i.e. mixed) bipartite entangled states. Here, distillation means that the target state is obtained with fidelity as close to 1 as desired as the size of the network increases when the pairwise entangled links support a constant amount of noise. While we prove an obstruction to multipartite distillation protocols after teleportation with channels with constant noise, we show that partial distillability is indeed possible if certain well-established graph-theoretic parameters that measure the connectivity in the network grow fast enough with its size. We give necessary as well as sufficient conditions for partial distillability in terms of these parameters and we moreover provide explicit constructions of networks with partial distillability and a relatively slow connectivity growth.
@@ II.1 Graph theory
In this subsection we provide the definition of all the graph-theoretic concepts that are used in this work. For a more detailed discussion we refer our readers to the well-established literature in the field such as [24, 25]. All graphs in this article are finite and simple. Thus, a graph is a pair $G=(V,E)$, where $V$ is a finite set of vertices and $E$ a set of edges, consisting of unordered pairs $(u,v)$ of vertices $u,v\in V$ ($u\neq v$). The order of $G$, denoted $|G|$, is the cardinality of the set of vertices $V$ and the size of $G$ is the cardinality of the set of edges $E$. Two vertices $u,v\in V$ are said to be adjacent if $(u,v)\in E$. We denote the set of edges adjacent to a vertex $v\in V$ by $E(v)$. The degree of a vertex $v\in V$, $\delta(v)=|E(v)|$, is the number of adjacent vertices to it, and we denote by $\delta_{\mathrm{min}}(G)=\min_{v\in V}\delta(v)$ ($\delta_{\mathrm{max}}(G)=\max_{v\in V}\delta(v)$) the minimal (maximal) degree of a graph $G$. A path is a sequence of edges joining subsequent vertices, $(v_{0},v_{1}),(v_{1},v_{2}),\dots,(v_{\ell-1},v_{\ell})$, and when $v_{j}\neq v_{i}$ for $i\neq j$ the path is said to be simple. A cycle is a path starting and ending on the same vertex. The length of a path is the number of edges in the path and the distance between two vertices $u,v\in V$, $d_{G}(u,v)$, is the length of the shortest path joining $u$ and $v$.
@@ II.2 Quantum (network) states and distillability
In this article we will always consider finite-dimensional Hilbert spaces. A quantum state is given by any positive semidefinite unit-trace operator $\rho$, referred to as density operator, acting on a given Hilbert space $\mathcal{H}$. The set of all density operators on $\mathcal{H}$ is denoted by $D(\mathcal{H})$. When a density operator is a rank-1 projector we say that the corresponding state is pure. When a quantum system is composed by $N$ subsystems, labelled by the elements of $N]:=\{1,2,\ldots,N\}$, the corresponding Hilbert space is given by $\mathcal{H}=\bigotimes_{k=1}^{N}\mathcal{H}_{k}$, where $\mathcal{H}_{k}$ is the Hilbert space corresponding to party $k$. In the bipartite case, $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$, we denote by
@@
$\rho(p)=p\phi^{+}+\frac{(1-p)}{d^{2}}\mbox{$1\hskip-2.84526pt{\bf l}$},$ (2)
where $p$ is referred to as the visibility and $1\hskip-2.84526pt{\bf l}$ stands for the identity operator on $\mathcal{H}$. It is known that the isotropic state is entangled if and only if (iff) $p>1/(d+1)$ [26] and it is a common noise model in which one produces a maximally entangled state with probability $p$ and white noise with probability 1p1-p.
@@
An $N$-partite quantum network state is described by a graph $G=(V,E)$ such that $|G|=N$. The vertices (i.e. the elements of $V$) represent the parties and the edges (i.e. the elements of $E$) that the corresponding pair of parties share some bipartite state to be specified. Thus, we denote the total Hilbert space by $\mathcal{H}_{G}$ and $\mathcal{H}_{G}=\bigotimes_{v\in V}\mathcal{H}_{v}$, where $\mathcal{H}_{v}$ represents the Hilbert space corresponding to the party represented by the vertex $v\in V$. Notice that there is a bijective identification between the elements of $V$ and those of $[N]$ and, thus, in a slight abuse of notation, we will refer to both vertices and parties as elements of these two sets equally. If vertices $u$ and $v$ share the state $\rho$, then we write $\rho_{e}$, where $e=(u,v)\in E$, and $\rho_{e}\in D(\mathcal{H}_{e})$, where $\mathcal{H}_{e}=\mathcal{H}_{ue}\otimes\mathcal{H}_{ve}$. Thus, $\mathcal{H}_{v}=\bigotimes_{e\in E(v)}\mathcal{H}_{ve}$. Furthermore, it also holds that $\mathcal{H}_{G}=\bigotimes_{e\in E}\mathcal{H}_{e}$, even though $\mathcal{H}_{e}$ does not correspond to the local Hilbert space of any party.
@@ II.3 Noisy teleportation
The main ingredient in the LOCC protocols that we are going to consider is teleportation [30, 31]. We denote by $\mathcal{T}:D(\mathcal{H}_{1}\otimes\mathcal{K}_{1}\otimes\mathcal{H}_{2})\to D(\mathcal{H}_{2})$ the channel implementing the standard bipartite teleportation protocol (where $\dim\mathcal{H}_{1}=\dim\mathcal{K}_{1}=\dim\mathcal{H}_{2}$ and $\mathcal{H}_{1}\otimes\mathcal{K}_{1}$ is the Hilbert space of one party and $\mathcal{H}_{2}$ is the Hilbert space of the other). That is, if $\phi^{+}\in D(\mathcal{K}_{1}\otimes\mathcal{H}_{2})$ denotes the maximally entangled state (cf. Eq. (1)), then
$\mathcal{T}(\tau\otimes\phi^{+})=\tau$ (7)
holds for every choice of input state $\tau\in D(\mathcal{H}_{1})$. We will be often using the entangled links in an isotropic network state to teleport quantum states through the network, which will be noisy when the visibility satisfies $p<1$. Applying the same protocol by using an isotropic state instead of a maximally entangled state yields
$\mathcal{T}(\tau\otimes\rho(p))=p\tau+(1-p)\frac{\mbox{$1\hskip-2.84526pt{\bf l}$}}{d}.$ (8)
We define the noisy teleportation channel $\mathcal{T}_{p}:D(\mathcal{H}_{1})\to D(\mathcal{H}_{2})$ as
$\mathcal{T}_{p}(\tau):=\mathcal{T}(\tau\otimes\rho(p)),$ (9)
@@
$\mathcal{T}_{p}(|i\rangle\langle j|)=p|i\rangle\langle j|+(1-p)\delta_{ij}\frac{\mbox{$1\hskip-2.84526pt{\bf l}$}}{d}.$ (10)
@@
Another option is to directly distribute a multipartite quantum state to some set of parties $V$ by means of a star graph $S=(V^{\ast},E)$, where $V^{*}=V\cup\{v_{0}\}$ and $v_{0}$ is its center. This additional party can locally prepare any state $\tau\in D(\mathcal{H}^{\prime}_{V})$ and teleport each share of it to each of the parties in $V$ by means of each edge in the star graph using the corresponding isotropic state as a resource for noisy teleportation. Thus, this corresponds to implementing the LOCC map from $D(\mathcal{H}_{S})$ to $D(\mathcal{H}^{\prime}_{V})$ given by $(\mathcal{T}_{p})^{\otimes|V|}(\tau)$ for any chosen $\tau\in D(\mathcal{H}^{\prime}_{V})$. If the parties in $V$ are linked to many additional parties giving rise to star graphs with different centers, they can iterate this process in order to end up with many noisy copies of the target state, to which they can apply multipartite distillation protocols as a final stage. We will in particular later on consider the paradigmatic GHZ $n$-qubit state
@@
$\sigma_{k}^{C}=\frac{1}{2}\left(\bigotimes_{\ell\notin C}|0\rangle_{\ell}\langle 0|+\bigotimes_{\ell\notin C}|1\rangle_{\ell}\langle 1|\right)\otimes\frac{\mbox{$1\hskip-2.84526pt{\bf l}$}_{C}}{2^{k}}.$ (15)
@@ III Limitation of multipartite distillation protocols
where #(j,k)\#(j,k) is the number of elements in $\{C\in\mathcal{C}(n,k)\mid\langle\Psi_{j}^{\pm}|\sigma_{k}^{C}|\Psi_{j}^{\pm}\rangle\neq 0\}$; that is, the number of non-vanishing terms in the sum $\sum_{C\in\mathcal{C}(n,k)}\langle\Psi_{j}^{\pm}|\sigma_{k}^{C}|\Psi_{j}^{\pm}\rangle$. Let $C\in\mathcal{C}(n,k)$, $j_{1}j_{2}\dots j_{n-1}$ the binary expression of $j$ and $j_{n}=0$ in order to use the same notation as above. Given $j$, Lemma 3 tells us #(j,k)\#(j,k) is the number of elements $C\in\mathcal{C}(n,k)$ such that for every $\ell_{1},\ell_{2}\notin C$ it holds $j_{\ell_{1}}=j_{\ell_{2}}$. For $i=0,1$, let us denote by $_{i}(j,k)$ the number of elements $C\in\mathcal{C}(n,k)$ such that $j_{\ell}=i$ implies $\ell\in C$. By definition, it holds $(j,k)=\#_{1}(j,k)+\#_{0}(j,k)$. It follows now that $_{1}(j,k)=0$ whenever $k<w_{H}(j_{1}j_{2}\dots j_{n-1}0)=w_{H}(j)$. Otherwise, $_{1}(j,k)$ must be the number of different subsets of $[n]$, with size $k$, containing the positions in which $j_{1}j_{2}\dots j_{n-1}0$ has ones. Since these $w_{H}(j)$ positions have to be assigned to $C$ always, $_{1}(j,k)$ is given by the different ways in which the remaining $n-w_{H}(j)$ positions can fill the remaining $k-w_{H}(j)$ slots in $C$, i.e.
$_{1}(j,k)=\binom{n-w_{H}(j)}{k-w_{H}(j)}$. (25)
On the other hand, for $_{0}(j,k)$ the $n-w_{H}(j)$ positions where $j_{1}j_{2}\dots j_{n-1}0$ has a zero must always be assigned to $C$ and, therefore, $_{0}(j,k)$ is given by the different ways in which the remaining $w_{H}(j)$ positions can fill the remaining $k-(n-w_{H}(j))$ slots in $C$, i.e.
$_{0}(j,k)=\binom{w_{H}(j)}{k+w_{H}(j)-n}$. (26)
@@ IV.1 A sufficient condition for partial distillability
If $(G_{n})_{n\in\mathbb{N}}$ is a graph sequence such that $\delta_{\mathrm{min}}(G_{n})$ is $\Omega(|G_{n}|)$ and $\lambda(G_{n})$ is $\omega(1)$ (i.e. $\lim_{n\to\infty}\lambda(G_{n})=\infty$), then it has partial distillability.
@@
and we assume that $M\geq 1$ since otherwise the claim is trivially true. For each $1\leq m\leq M$, we will construct the spider graph $S_{m}$ out of $|V_{0}|-1$ paths $P_{m,k}$, $1\leq k\leq|V_{0}|-1$ joining $v$ with $v_{k}$. To find these paths we use an inductive construction based on a family of subgraphs $G_{m,k}$, with $G_{1,1}=G$. If $G_{m,k}$ is connected, we define $P_{m,k}$ to be the shortest path inside $G_{m,k}$ joining $v$ with $v_{k}$. We then define $G_{m,k+1}$ to be the subgraph of $G_{m,k}$ obtained by removing all the edges in $P_{m,k}$. Once the last leg of the spider is obtained (i.e., the path $P_{m,|V_{0}|-1}$), we define $S_{m}$ to be the union of every path in $\{P_{m,k}\}_{k=1}^{|V_{0}|-1}$ and start building the next spider graph, starting over from $G_{m+1,1}=G_{m,|V_{0}|}$. Notice then that by construction the spider graphs are guaranteed to be edge-disjoint.
For the previous construction to work and to prove the claim, we need to ensure for all $1\leq m\leq M$ and all $1\leq k\leq|V_{0}|-1$ that $G_{m,k}$ is connected and that the length of the path $P_{m,k}$ is upper bounded by $5/c$. We know that $G_{1,1}=G$ is connected by hypothesis, and, by Eq. (34), that $\operatorname{diam}(G_{1,1})\leq 3/c<5/c$, so that $P_{1,1}$ fulfills the desired bound on its length. Thus, we can use induction over $j$, where we use the bijective mapping $j(k,m):=(m-1)|V_{0}|+k$, assuming that $1\leq j\leq M|V_{0}|-1$ (so that $1\leq m\leq M$ and $1\leq k\leq|V_{0}|-1$). We have just verified that the first element in this set satisfies our claim, so we now assume that $G_{m,k}$ is connected and that the length of $P_{m,k}$ is upper bounded by $5/c$ and we need to show that the same holds for the graph indexed by $j(k,m)+1$ as long as $j\leq M|V_{0}|-2$. Note that $G_{m,k}$ is built by removing a simple path from $G_{m,k-1}$ (or $G_{m-1,|V_{0}|-1}$ if $k=1$), and therefore in each step the minimal degree of $G_{m,k}$ can at most be decreased by 2 from the previous step, which implies
@@
We will explicitly build a protocol that achieves partial distillation under the given premises. Since $\delta_{\mathrm{min}}(G)\leq|G|-1$ for any graph $G$, $\delta_{\mathrm{min}}(G_{n})$ being $\Omega(|G_{n}|)$ implies that there exists $0<c<1$ and $n_{1}>0$ such that $\delta_{\mathrm{min}}(G_{n})\geq c|G_{n}|$ for every $n>n_{1}$. In the remainder of the proof we assume that $n$ is large enough so that the above inequality holds. Fix